Math Destroyer question

Started by DK17
This forum made possible through the generous support of SDN members, donors, and sponsors. Thank you.
Get help with your application

Use all the free resources available to you from SDN: articles, guides, expert advising, forums discussions, and school research.

DK17

New Member
5+ Year Member
Advertisement - Members don't see this ad
Fellow SDN'ers, I need clarity on this question.

Half the population of Green Village is vegetarians, and 35% ride bicycles. If 20% are both vegetarian and ride bicycles, what is the probability that a person chosen at random from Green Village is either a vegetarian or rides a bicycle.

A. 95%
B. 65%
C. 12%
D. 83%
E. 24%

Solution in the back says its B, but if I think of it, why is it simply not: 50/100 + 35/100 = 85/100 or 85% either a vegetarian or rides a bicycle? Why does the 20% even matter? Sorry if this is a stupid question, just been bothering me and lingering in the back of my mind, need an answer -_-.
 
Since 25% of people do both, then that's an overlap that needs to be removed or else you're counting the same people twice. (50-20) + 35 (or (35-20) +50) would be a better way of looking at it.
This is a problem where you don't even have to try to reason it out and possibly confuse yourself. I know a lot of people have trouble with the ven diagrams. If you can just remember 1 or 2 formulas then you'll be set.

Probability Rules - this link details some of them
 
Ahhhh, better understood by, they want either/or, not including both. You guys are awesome, it seems like I didn't read the question thoroughly enough. Thanks for clearing that up guys [: